# code-for-first-row = , code-for-last-row = , code-for-first-col = , code-for-last-col =

### v1

@@ 4 Clustering Partially Ordered Data
Cluster analysis is a multivariate data analysis technique for identifying structural patterns and relationships within complex datasets with the goal of grouping similar data points into clusters. In the process, it attempts to discover the internal structure of a data set by considering the similarity between objects based on some type of attributes that the objects may possess in varying degrees. The ordinal model for clustering with posets was introduced in [116], where it is shown that the characterization of flat cluster methods[115] leads to a universal mapping problem in the theory of partially ordered sets. A persistent underlying theme involves generalized notions of adjoints of order preserving mappings between posets [117]. In addition, [117] attempts to develop a clustering scheme based on dissimilarities measured in posets. Many of the cluster methods that are used in the construction of classificatory systems operate on data in the form of a dissimilarity coefficient on a set of objects [21]. There are two methods in cluster analysis, namely hierarchical and non-hierarchical. Mathematically, hierarchical clustering involves the creation of a nested sequence of partitions of a set, whereas non hierarchical clustering deals with a single partition.The hierarchical clustering method consists of the agglomerative and divisive methods. Divisive clustering use a top-down approach to combine all the data points as a single cluster and then divide them as the distance between them increases. On the other hand, agglomerative clustering is a bottom-up approach which divides the data points into different clusters and then aggregates them as the distance decreases. Agglomerative clustering can be categorized into single linkage and complet linkage. Single linkage is a method that focuses on minimum distances or nearest neighbor between clusters while complete linkage focuses on maximum distance or furthest neighbor between clusters [285]. Sabara et al. [190] use the single linkage and complete linkage agglomerative hierarchical clustering method on the Hasse graph of posets in a multidimensional data analytics setting. Similarly, [189, 122] also develop methods that combine poset methodology with hierarchical clustering from a multidimensional data perspective. In particular, [189] develops a method of hierarchical stratification which is interpretable using the partial order Hasse graph and the evaluation process is such that it takes into consideration the characteristics of multi-dimensional indicators, division into several layers whereby different layers have different properties and characteristics.
@@
Janowitz [118] presents a comprehensive review of clustering including through the use of lattices that generalize trees. Ontologies represent data relationships as hierarchies of possibly overlapping classes. Ontologies are closely related to clustering hierarchies.Dissimilarity is a common intermediary used by clustering methods to classify data. In [131], it is shown that modeling ontologies as posets over the subset relation enables classical clustering algorithms which take dissimilarity matrices as inputs always incorporate all available information, as such there is no loss of information. This served as the basis for the development of a clustering algorithm, that generates a partially ordered set of clusters from a dissimilarity matrix. Dissimilarity matrices simplify some of the problems associated with clustering high-dimensional datasets, since their size is only a function of the number of objects $\mathcal{O}(N^{2})$ , and independent of the objects’ dimensions.
@@
Formal Concept Analysis (FCA) is an approach for identifying structure present in domains. The FCA can be seen as a conceptual clustering method. Michalski(1980)[321] introduced conceptual clustering as a new branch of machine learning. Since its introduction there has been significant interest in the different machine learning tasks that can utilize conceptual clustering methods. Work in Artificial Intelligence has concentrated significantly on developing mechanisms for classification, and the conceptual representations has been pivotal in supporting this effort. Conceptual clustering [69] is concerned with the problem of grouping unlabelled objects into classes. Three key features of conceptual clustering methods are:
@@
    Output classes (often referred to as concepts) are arranged into a hierarchy based on their generality/specificity
@@
    the process of class formation is incremental; that is, the processing of the $n$th object does not require extensive reprocessing of the $(n-1)$ objects already seen. description
@@
The theory of concept or galois lattices provides a simple and formal approach to conceptual clustering. Carpineto[68] develop the clustering algorithm GALOIS which is able to determine the concept lattice corresponding to a given set of objects. It was demonstrated to be useful for class discovery and class prediction with a the time complexity of each update ranging from $O(n)$ to $O(n^{2})$ where n is the number of concepts in the lattice. Fisher [322]presented conceptual clustering as an extension of numerical taxonomy. In [240], the theoretical background of the application of Hierarchical Cluster Analysis HCA to improve on the understanding of Hasse Diagrams derived from the ranking process is presented. The use of HCA is based on the idea of reducing the number of elements in a poset by the selection of representatives from the original set used for ranking. As such the clusters derived from HCA can be interpreted as similarity classes in which one of its members (the nearest to the centre of the cluster) can be selected as representative of its class. Generally, incremental hierarchical conceptual clustering algorithms work by defining a set of hierarchy-change operators and carrying out a hill-climbing search over the space of possible concept hierarchies with the goal of determining a “good” hierarchy [214]. These algorithms have been implemented by several incremental hierarchical conceptual clustering systems have been developed over the years([199, 200, 201, 202, 203, 204, 210]). Although they differ in many respects (e.g., concept representation, hierarchy evaluation criteria, disjoint versus nondisjoint classes), they usually solve the problem by defining a set of hierarchy-change operators and carrying out a hill-climbing search through the space of possible concept hierarchies aimed at finding a “good” hierarchy (Gennari et al., 1989). These systems perform well in many experimental settings, but the criteria they use to modify a [69]. Another major developemnt in hierarchical conceptual clustering is its evaluation metrics. Jonyer et al. [213, 212] provides an objective evaluation metric that reflects the major requirements and tradeoffs of a good quality clustering. Yoneda et al.[282] present a method that can learn a graph structured representation from multivariate data where each node represents a cluster of data points and each edge represents the subset-superset relationship between clusters. The main idea of the method use formal concept analysis (FCA), which can extract hierarchical relationships between clusters based on the algebraic closedness property which leads to a directed graph representation.
@@ 5 Multidimensional Data Analysis
Multidimensional data analysis[249, 271, 248] allows the observation of data from various viewpoints.These data can most often be organized into meaningful hierarchies [272]. Multi-indicator models [250] are vey useful in the domain of multidimensional data analysis. Multiple indicators consist of two or more ”alternative” measures of the same concept. That is, a causal model contains alternative measures of the same thing. For example., two different ways of asking how satisfied you are with a service. Most of its familiar applications can be found in fields that involve the evaluation of subjective data. Many public opinion surveys employ this kind of data, usually measured among many thousands of individuals [49, 114]. Due to the complexity of multi-indicator systems[46], quantifying these indicators can be difficult and time expensive. Specifically, when dealing with nominal or ordinal multivariate data, a preliminary step is often scaling them with quantitative data. However, this can lead to inconsistent results[85] The most popular approach for quantifying multi-indicator data values is based on weighted-average or other aggregation techniques(additive[273], hierarchical[274] etc.) which is then used in a ranking procedure [39]. Ranking data represent a peculiar form of multivariate ordinal data taking values in the set of permutations [CITATION:TITLE:Mollica, PLMIX:AnRpackageformodelingandclusteringpartiallyrankeddata]. The mathematical aggregation of a set of individual indicators that measure multi-dimensional concepts leads to a composite indicator [275]. In the field of social sciences, composite indicators based on scoring or index functions are a commonly used method for ranking or measuring objects based on different properties[17]. A composite indicator may include several dimensions, where the dimensions represent different domains or aspects of the phenomenon being measured. For example, a composite indicator of well- being may cover dimensions such as income, employment, health and education. The problem of ranking subjects based on some ordinal scale is a fundamental issue in economics and social statistics. In addition, the validity and robustness of the results obtained from composite indices or other multidimensional data sources through aggregation has been challenged particularly due to the unavoidable aspect of subjectivity in their calculations. In addition, aggregative methods tend to oversimplify the complexity of a phenomenon and they also produce the same results for different situations [3, 5]. Moreover, such data often lack clear ordering criteria and many ambiguities arise when trying to compare individuals consistently[85]. In order to address these issues, many studies ([17, 31, 142, 87, 46, 49, 96, 10, 14, 48, 91, 61] etc. ) advocate the use of partially ordered set theory, which requires subjectivity only in the choice of the properties to be considered. The key motivation for advocating a poset oriented approach derives from the observation that classical aggregative methods of analysis may not clearly express the true complex nature of certain phenomena. Partial orders help to reveal why an object of interest holds a certain ranking position and how much it is subject to change if a composite indicator is upgraded [46]. This methodology is convenient in providing comparable indices or ranks of objects while reducing subjectivity to a minimum. The main advantages of managing multivariate data for ordering purposes based on the notion of posets for the purpose of anking are summarized as follows.
@@
  • Partial order theory and the related Hasse Diagram Technique provides tools for dealing with multidimensional systems of ordinal data which avoids the procedure of performing variable aggregation into composite indicators. As such, there is no necessity for converting ordinal scores into numerical values which can lead to inconsistency in the evaluation of the real phenomena under consideration [87, 10].
@@
  • The evaluative process for ranking using partial order theory is advantageous in that the comparison of the objects of interest is done simultaneously for all indicators, without the need of any prior aggregation [63].
@@
  • Partial order ranking offers a non-parametric method that neither includes any assumptions about linearity nor any assumptions about distribution properties [53].
@@
  • Ranking by Hasse Diagram Technique has the advantage that it can be performed without any normative constraints. The theory of partially ordered sets and its graphical representation (Hasse diagrams) does not need any additional information to sort objects. Beyond sorting, many conclusions can be drawn from Hasse diagrams because they represent a well-defined mathematical structure [31, 27].
@@
  • Posets are used to define the structure of comparabilities underlying multi-indicator systems [5]. Once the structure is defined, we can analyse it by using mathematical tools. The first goal of this analysis is the evaluation of scores to reduce the dimensionality of the phenomenon. We deal with quantitative variables so we analyse the resulting structure of comparabilities (the poset), without evaluating any difference or ratio to their values to get distances or proportions. Therefore, there is no loss of information.
@@
We summarize how to obtain poset structures from subjective data set as follows. Let $k$ be the number of subjective indicators related to some observed phenomenon and let $n$ denote the number of individuals in the study such as a survey. Furthermore, let $\text{A}=\{q_{1},q_{2}...,q_{k}\}$ be such that each element of the set $A$ represents a different variable corresponding to a subjective indicator. Score the response information of each subjective indicator on an $m$ degree ordinal scale $(y_{1},y_{2}...,y_{m})$ which are ordered in such a way that $y_{1}<y_{2}<...<y_{m}.$ For each statistical unit associate the set of degrees on the variables. Each of the $n$ individual profiles is determined by the $m$ ordinal scores given in the set $B$ on each of the $k$ indicators in the set $A$. This is followed by pairwise comparison of all the $n$ individuals using their individual profiles corresponding to the ordinal scale in $B.$This evaluation procedure if not totally ordered leads to the poset structure which can be depicted by a Hasse graph of size $n.$ For instance, using this set up for a four degree scale, the profile $(y_{2},y_{3},y_{2},y_{3})$ is better than $(y_{4},y_{3},y_{3},y_{1}),$ but $(y_{3},y_{3},y_{4},y_{2})$ and $(y_{4},y_{3},y_{3},y_{2})$ are incomparable. Different types of methodology has been put forward which use poset theory for the analysis of multivariate datasets that have attributes which require multidimensional approach. Some of these are based on 1) fuzzy set approach [86, 85], partial order linear extension [87] 3) average rank [49] 4) poset theory with Adjusted Mazziotta–Pareto Index [4] 5) interpersonal comparability [155] 6) fuzzy-first order dominance [96] and so forth.
@@
  • The fuzzy set approach was originally developed in the seminal work of Zadeh (1965)[251] as a class of sets that allows its members to have different degree of membership It can address uncertainty associated with vague concepts by allowing partial membership to a set. Fuzzy set approach has been used for the statistical evaluation of ordinal data associated to socio-economic phenomena in combination with posets in order to overcome the problems associated with the classical aggregative approach based on composite indicators[86, 85, 12]. The main feature of this approach is that the measurement level of the data is fully respected, avoiding any kind of not proper scaling.
@@
  • In [87] a ranking method is developed based on the linear extensions from partial orders. Particulaly, it aims to solve the problem of ranking a finite collection of objects which can be represented as a cloud of points having a suite of indicator values corresponding to each member of the collection. It solves the problem of arbitrary assignment and aggregation of composite numerical score for each object , by viewing the relative positions as defining a partial ordering and any given pair of objects may not necessarily be comparable. The interval having the possible ranks assigned to each object can then be determined from the Hasse diagrams of the partial orders corresponding to the collection of all rankings that are compatible with the linear extensions from partial orders. In cases involving very large datasets which requires the enumeration a large number of linear extensions that can result to computational intractability, it is suggetsed in [142] to use the method of discrete Markov chain Monte Carlo for such intractable problems
@@
  • Caperna & Boccuzzo[49] suggest a methodology for dealing with big complex data when some of the attributes require the measurement of complex concepts on ordinal and/or dichotomous scales. The method of average rank as a synthetic measure of a complex concept was motivated by the notion that poset theory is of paramount importance for dealing with ordinal data avoiding the use of scaling procedures. The main idea consist of a procedure for sampling units from a big population using a simple criterion to summarize the resulting values appropriately. This is followed by an application of the central limit theorem which makes it possible to compare the results obtained from different groups using statistical tests on the means. The Height of Groups by Sampling method is then used to compare the average rank among groups that are defined by one or more socio-demographic variables influencing the level of the measurable complex concept.
@@
  • In [143] some procedures for n extracting insights on ordering properties that are embodied in multivariate datasets and applicable in configuring sets of indicators are reviewed.
@@
  • The applied methodology [4] for synthesizing the evolution of multi-indicator systems over time is based on the partial order theory and the aggregative method of the Adjusted Mazziotta–Pareto Index, to one of the 15 sustainable development goals. In particular, the role of posets was to defne the structure of comparabilities underlying multi-indicators systems. This is then followed by a mathematical evaluation of scores to reduce the dimensionality of the phenomenon. Temporal posets are then obtained by merging the diferent posets. Furthermore, Aan embedded five level scale divided into( minimum, maximum, first, second and theird quartiles of the indicator values are also used to improve the quality of measurement.
@@
  • Sen [155] propose a framework for interpersonal comparability in the context of aggregating the welfare measures of individuals. An aggregation relation between any pair of social states which is a quasi-ordering and identical with respect to the Pareto quasi-ordering under noncomparability, and is a complete ordering under unit comparability, and under full comparability are all considered.
@@
  • Fattore[92]: proposed a procedure for the assessment of multidimensional deprivation with ordinal data, based on partial order theory. The procedure focuses on achievement profiles and their comparabilities and incomparabilities The evaluation problem is treated in terms of multidimensional comparisons among profiles. The usefulness of the procedure is noted in its consideration for both vagueness and intensity of multidimensional deprivation and produces a complete set of synthetic indicators for the description of deprivation on statistical populations. In addition, t also allows for the introduction of information on attribute importance and it facilitates the handling missing data.
  • In [96] the procedure ”fuzzy-first order dominance ”(F-FOD), employs concepts and tools from partially ordered set theory and from fuzzy relational calculus with the goal of overcoming the main limitations of previously developed algorithms for FOD analysis. F-FOD is useful in that it produces full pairwise comparison matrices which allows for partial orderings and rankings of the statistical units to be derived from the input data.
The sorting approach using the Hasse Diagram Technique (HDT)[32], can be summarized as follows. Consider a dataset of objects to be sorted with respect to some criteria or property. Let m(i) be an appropriate attributes by which the objects are ordered. Furthermore, let $x$ and $y$ represent any two objects with $m(i,x)$ the value of the $i^{th}$ attribute of the object $x$($i=1,..,n$).Then $x\geq y$(which can mean that $x$ is evaluated better than or equal to $y$) if and only if $m(i,x)\geq m(i,y)$ for every $i=1,...,.,n$

### v2

@@ 4 Clustering Partially Ordered Data
Cluster analysis is a multivariate data analysis technique for identifying structural patterns and relationships within complex datasets with the goal of grouping similar data points into clusters. In the process, it attempts to discover the internal structure of a data set by considering the similarity between objects based on some type of attributes that the objects may possess in varying degrees. The ordinal model for clustering with posets was introduced in [174], where it is shown that the characterization of flat cluster methods[173] leads to a universal mapping problem in the theory of partially ordered sets. A persistent underlying theme involves generalized notions of adjoints of order preserving mappings between posets [175]. In addition, Janowitz[175] attempts to develop a clustering scheme based on dissimilarities measured in posets. Many of the cluster methods that are used in the construction of classificatory systems operate on data in the form of a dissimilarity coefficient on a set of objects [35]. There are two methods in cluster analysis, namely hierarchical and non-hierarchical. Mathematically, hierarchical clustering involves the creation of a nested sequence of partitions of a set, whereas non hierarchical clustering deals with a single partition.The hierarchical clustering method consists of the agglomerative and divisive methods. Divisive clustering uses a top-down approach to combine all the data points as a single cluster and then divides them as the distance between them increases. On the other hand, agglomerative clustering is a bottom-up approach which divides the data points into different clusters and then aggregates them as the distance decreases. Agglomerative clustering can be categorized into single linkage and complete linkage. Single linkage is a method that focuses on minimum distances or nearest neighbor between clusters while complete linkage focuses on maximum distance or furthest neighbor between clusters [240]. Sabara et al.[251] use the single linkage and complete linkage agglomerative hierarchical clustering method on the Hasse graph of posets in a multidimensional data analysis setting. Similarly, Wu et al.[300] and Kardaetz et al.[185] also develop methods that combine poset methodology with hierarchical clustering from a multidimensional data perspective. In particular, Wu et al.[300] develop a method of hierarchical stratification which is interpretable using the partial order Hasse graph, and the evaluation process is such that it takes into consideration the characteristics of multi-dimensional indicators, division into several layers whereby different layers have different properties and characteristics. Janowitz [176] presents a comprehensive review of clustering including through the use of lattices that generalize trees.
@@
Ontologies represent data relationships as hierarchies of possibly overlapping classes. Ontologies are closely related to clustering hierarchies. Dissimilarity is a common intermediary used by clustering methods to classify data. In Liu et al.[209], it is shown that modeling ontologies as posets over the subset relation enables classical clustering algorithms which takes dissimilarity matrices as inputs and as such it incorporates all available information and therefore there is no loss of information. This served as the basis for the development of a clustering algorithm that generates a partially ordered set of clusters from a dissimilarity matrix. Dissimilarity matrices simplify some of the problems associated with clustering high-dimensional datasets, since their size is only a function of the number of objects $\mathcal{O}(N^{2})$, and independent of the object’s dimensions.
@@
Formal concept analysis(FCA) is a powerful tool for investigating and identifying hidden structures in large data sets. Janowitz[176] introduces a mathematical treatment of the subject ordinal and relational clustering using FCA. The FCA can be seen as a conceptual clustering method. Michalski[220] introduced conceptual clustering as a new branch of machine learning. Since its introduction there has been significant interest in the potential machine learning tasks that can utilize conceptual clustering methods. Work in artificial intelligence has concentrated significantly on developing methods for classification, and the conceptual representations has been pivotal in supporting this effort. Conceptual clustering[90] is concerned with the problem of grouping unlabelled objects into classes. Furthermore, Carpineto[90] presents the three key features of conceptual clustering methods as follows:
@@
    Output classes (also referred to as concepts) are arranged into a hierarchy based on their generality or specificity;
@@
    The process of class formation is incremental in the sense that the processing of the $n$th object does not require extensive reprocessing of the previously processed $(n-1)$th objects.
@@
The theory of concept or Galois lattices provides a simple and formal approach to conceptual clustering. The Galois (or concept) lattice which is generated from a binary relation has been proven to be useful for many applications. The construction of Galois lattice [146] can be considered a conceptual clustering method because it results in a concept hierarchy. Carpineto[89] developed the clustering algorithm GALOIS which is able to determine the concept lattice corresponding to a given set of objects. It was demonstrated to be useful for class discovery and class prediction, with the time complexity of each update ranging from $O(n)$ to $O(n^{2})$ where n is the number of concepts in the lattice. Fisher[133] presented conceptual clustering as an extension of numerical taxonomy. In Restrepo et al.[242] the theoretical background of the application of Hierarchical Cluster Analysis( HCA) to improve on the interpretability of Hasse Diagrams derived from the ranking process is presented. The use of HCA is based on the idea of reducing the number of elements in a poset by the selection of representatives from the original set used for ranking. As such the clusters derived from HCA can be interpreted as similarity classes in which one of its members (the nearest to the centre of the cluster) can be selected as representative of its class. The process of determining the concepts in formal concept context is time-consuming, as it is considered an NP problem. Zhang et al. [309] present a hierarchical conceptual clustering analysis based on distance function for the purpose of determining the concepts in formal context. In this way, the discovery of all concepts is facilitated by subsetting of the feature set used in the clusters. On the other hand Markov[214] uses hierarchical clustering based on a generalization operator to induce a lattice structure to the clusters. As such it maximizes the overall clustering quality since the local evaluation goes through all levels of the hierarchy in a bottom-up fashion. Yoneda et al.[305] present a method that can learn a graph structured representation from multivariate data where each node represents a cluster of data points and each edge represents the subset-superset relationship between clusters. The main idea of the approach uses formal concept analysis to extract hierarchical relationships between clusters based on the algebraic closedness property which leads to a directed graph representation. Similarly, other graph-based hierarchical conceptual clustering methods which can be applicable to partially ordered datasets have been studied in [180, 181]. In clustering, the frame of discernment is the set of all partitions of a finite set $E$, denoted $\mathcal{P}(E)$ This set can be partially ordered using the following relation. A partition p is said to be finer than a partition $p^{\prime}$(or, equivalently $p^{\prime}$ is coarser than p) if the clusters of p can be obtained by splitting those of $p^{\prime};$ therefore $p\preccurlyeq p^{\prime}.$ In this case the poset $\left(\mathcal{P}(E),\preccurlyeq\right)$ can be obtained. Ensemble clustering aims at combining the outputs of several clustering algorithms (clusterers) to form a single clustering structure. Using evidential reasoning based on the Dempster-Shafer theory[109] by assuming that there exists a “true” partition $p^{*}$ and that each clusterer provides evidence about $p^{*}.$ The evidence from multiple clusterers can then be combined to draw plausible conclusions about $p^{*}$ in situations which are not immediately evident.
@@ 5 Multidimensional Data Analysis
Multidimensional data analysis was primarily established as a consequence of the development of relational databases and OLAP(On-Line Analysis Processing) by E.. Codd in 1993. The main idea of OLAP is based on the notion of dimensions which provides a flexible approach to interpreting data from several angles. That is, multidimensional data analysis[236, 161, 139, 206] allows the observation of data from multiple viewpoints. In statistics, econometrics and related fields, multidimensional analysis (MDA) is a data analysis process that groups data into two categories, namely data dimensions and measurements. Dimensions are a set of axis on which analysis is carried out. Measurements indicate values which are associated with different dimensional entities. As such the analyses and categorization of data based on multiple dimensions and measures constitutes multidimensional data analysis. In MDA, data analysis is based on hierarchy, sequential property and dependency relationships in the dimensions [161]. This form of data can most often be organized into meaningful hierarchies [42, 236]. Multi-indicator models[269] are very useful in the domain of multidimensional data analysis. Multiple indicators consist of two or more “alternative” measures of the same concept. That is, a causal model contains alternative measures of the same thing. For example, two different ways of asking how satisfied an individual is with a service. Most of its familiar applications can be found in fields that involve the evaluation of subjective data. Many public opinion surveys employ this kind of data, usually measured among many thousands of individuals [71, 168]. Many environmental systems and infrastructure systems [227] are monitored using a set of indicator values that assess several features of site condition. Peculiarly, the values of the different indicators often convey different comparative messages for the study sites. It is also useful in the mathematical modelling of business processes, where large and complex data is involved across multiple geographic regions, channels and products. Due to the complexity of multi-indicator systems[57, 65], quantifying these indicators can be difficult and time expensive. Specifically, when dealing with nominal or ordinal multivariate data, a preliminary step is often scaling them with quantitative data. However, this can lead to inconsistent results[118]. The most popular approach for quantifying multi-indicator data values is based on weighted-average or other aggregation techniques(additive[169], hierarchical[232] etc.) which is then used in a ranking procedure[59]. Ranking data represents a peculiar form of multivariate ordinal data taking values in the set of permutations[224]. The mathematical aggregation of a set of individual indicators that measures multi-dimensional concepts leads to a composite indicator [226]. In the field of social sciences, composite indicators based on scoring or index functions are a commonly used method for ranking or measuring objects based on different properties[22]. A composite indicator may include several dimensions, where the dimensions represent different domains or aspects of the phenomenon being measured. For example, a composite indicator of well- being may cover dimensions such as income, employment, health and education. The problem of ranking objects based on some ordinal scale is a fundamental issue in economics and social statistics. In addition, the validity and robustness of the results obtained from composite indices or other multidimensional data sources through aggregation has been challenged particularly due to the unavoidable aspect of subjectivity in their calculations. In addition, aggregative methods tend to oversimplify the complexity of a phenomenon and they also produce the same results for different situations [3, 2]. Moreover, such data often lacks clear ordering criteria and many ambiguities arise when trying to compare individuals consistently[118]. In order to address these issues, many studies ([22, 50, 233, 120, 57, 71, 129, 12, 18, 69, 124, 82] etc.) advocate the use of partially ordered set theory, which requires subjectivity only in the choice of the properties to be considered. The key motivation for advocating a poset oriented approach derives from the observation that classical aggregative methods of analysis may not clearly express the true complex nature of certain phenomena. Partial orders help to reveal why an object of interest holds a certain ranking position and how much it is subject to change if a composite indicator is upgraded [57]. This methodology is convenient in providing comparable indices or ranks of objects while reducing subjectivity to a minimum. The main advantages of managing multivariate data for ordering purposes based on the concept of posets and its properties are summarized as follows.
@@
  • Partial order theory provides tools for dealing with multidimensional systems of ordinal data which avoids the procedure of performing variable aggregation into composite indicators. As such, there is no necessity for converting ordinal scores into numerical values which can lead to inconsistency in the evaluation of the real phenomena under consideration [120, 12].
@@
  • The evaluative process for ranking using partial order theory is advantageous in that the comparison of the objects of interest is done simultaneously for all indicators without the need of any prior aggregation [84].
@@
  • Partial order ranking offers a non-parametric method that neither includes any assumptions about linearity nor any assumptions about distribution properties [74]. That is, it is parameter free and as such there are no assumptions needed about the statistical distributions of the attributes.
@@
  • The theory of partially ordered sets and its graphical representation in the form of Hasse graphs does not need any additional information to sort objects. Beyond sorting, many conclusions can be drawn from Hasse diagrams because they represent a well-defined mathematical structure [50, 45].
@@
  • Posets are used to define the structure of comparabilities underlying multi-indicator systems [2]. The poset structure permits the analyses of the phenomenon under study using mathematical tools. A key goal of this analysis is the evaluation of scores to reduce the dimensionality of the phenomenon. In this case, quantitative variables are used for the analysis since the resulting structure of comparabilities defines the poset. Thus, there is no loss of information since it does not involve the possible evaluations of any difference or ratio of their values to obtain distances or proportions.
  • Partial order recognizes that the values of different indicators often convey a different comparative message and most often there is no unique way of solving many ranking problems [227], while taking all of the indicator information into account. For example, item A may be considered better than item B with respect to the first indicator, but item B is considered superior to item A with respect to the second indicator. The different comparative messages of an indicator set complicates the process of prioritization of the items based on the indicator scores. In particular, the properties of partially-ordered sets and rank matrices can provide additional useful information to aid decision making beyond an overall ordering of the items.
  • The presentation by a Hasse diagram of posets avoids the arbitrariness in constructing a ranking index [50].
  • Hasse diagrams of posets allows a holistic view on all the objects which are to be ranked, without introducing artificial ranking indices [46].
@@
We summarize how to obtain poset structures from subjective data set as follows. Let $k$ be the number of subjective indicators for the study of some observed phenomenon and let $n$ denote the number of individuals in the study. Furthermore, let $\text{A}=\{q_{1},q_{2}...,q_{k}\}$ be such that each element of the set $A$ represents a different variable corresponding to a subjective indicator. Score the response information of each subjective indicator on an $m$ degree ordinal scale $(y_{1},y_{2}...,y_{m})$ which are ordered in such a way that $y_{1}<y_{2}<...<y_{m}.$ For each statistical unit associate the set of degrees to the variables. Each of the $n$ individual profiles is determined by the $m$ ordinal scores given in a set $B$ on each of the $k$ indicators in the set $A$. This is followed by pairwise comparison of all the $n$ individuals using their individual profiles corresponding to the ordinal scale in $B.$ This evaluation procedure if not totally ordered leads to a poset structure which can be depicted by a Hasse graph of size $n.$ For instance, using this set up for a four degree scale, the profile $(y_{2},y_{3},y_{2},y_{3})$ is better than $(y_{4},y_{1},y_{3},y_{1}),$ but $(y_{3},y_{3},y_{4},y_{2})$ and $(y_{4},y_{3},y_{3},y_{2})$ are incomparable. Different types of methodology have been put forward which use poset theory for the analysis of multivariate datasets that have attributes which require a multidimensional approach. Some of these are based on 1) fuzzy set approach [119, 118], partial order linear extensions [120] 3) average rank [71] 4) poset theory with Adjusted Mazziotta–Pareto Index [1] 5) interpersonal comparability [258] 6) fuzzy-first order dominance [129] and so forth. These approaches are summarized as follows.
@@
  • The fuzzy set approach was originally developed in the seminal work of Zadeh (1965)[307] as a class of sets that allows its members to have different degrees of membership. It can address uncertainty associated with vague concepts by allowing partial membership to a set. Fuzzy set approach has been used for the statistical evaluation of ordinal data associated to socio-economic phenomena in combination with posets, in order to overcome the problems associated with the classical aggregative approach based on composite indicators[119, 118, 16]. The main feature of this approach is that the measurement level of the data is fully respected, avoiding any form of improper scaling.
@@
  • In Fattore et al.[120] a ranking method is developed based on the linear extensions from partial orders. Particularly, it aims to solve the problem of ranking a finite collection of objects which can be represented as a cloud of points having a suite of indicator values corresponding to each member of the collection. It solves the problem of arbitrary assignment and aggregation of composite numerical score for each object by viewing the relative positions as defining a partial ordering and any given pair of objects may not necessarily be comparable. The interval having the possible ranks assigned to each object can then be determined from the Hasse diagrams of the partial orders corresponding to the collection of all rankings that are compatible with the linear extensions from partial orders. In cases involving very large datasets which require enumerating a large number of linear extensions which can result to computational intractability, it is suggested in [233] to use the method of Discrete Markov Chain Monte Carlo(MCMC) for such problems. The MCMC method involves the sampling of linear extensions[68], followed by the computation of cumulative frequency ranking based on the sampled linear extensions. In Lerche[140], the method of calculating ranking probabilities based on random linear extensions is highlighted. The approach based on random linear extensions can be very useful for the prediction of ranking probabilities and computation of the average rank for larger partially ordered sets, which are practically impossible to handle using a very large set of generated linear extensions.
@@
  • Caperna & Boccuzzo[71] suggest a methodology for dealing with big complex data when some of the attributes require the measurement of complex concepts on ordinal or dichotomous scales. The main idea consist of a procedure for sampling units from a big population using a simple criterion to summarize the resulting values appropriately. This is followed by an application of the central limit theorem which makes it possible to compare the results obtained from different groups using statistical tests on the means. The Height of Groups by Sampling method is then used to compare the average rank among groups that are defined by one or more socio-demographic variables influencing the level of the measurable complex concept.
@@
  • In Patil et al. [234], a procedures for extracting insights on ordering properties that are embodied in multivariate datasets and applicable in configuring sets of indicators are reviewed.
@@
  • The applied methodology described in Alaimo et al[1] for synthesizing the evolution of multi-indicator systems over time is based on partial order theory and the aggregative method of the Adjusted Mazziotta–Pareto Index, to one of the 15 sustainable development goals. In particular, the role of posets was to define the structure of comparabilities underlying multi-indicators systems. This is then followed by a mathematical evaluation of scores to reduce the dimensionality of the phenomenon. Temporal posets are then obtained by merging the different posets. Furthermore, an embedded five level scale divided into( minimum, maximum, first, second and third quartiles of the indicator values are also used to improve the quality of measurement.
@@
  • Sen [258] proposes a framework for interpersonal comparability in the context of aggregating the welfare measures of individuals. An aggregation relation between any pair of social states which is a quasi-ordering and identical with respect to the Pareto quasi-ordering under noncomparability, and is a complete ordering under unit comparability, and under full comparability are all considered.
@@
  • Fattore[125] proposes a procedure for the assessment of multidimensional deprivation with ordinal data, based on partial order theory. The procedure focuses on achievement profiles and their comparabilities and incomparabilities The evaluation problem is treated in terms of multidimensional comparisons among profiles. The usefulness of the procedure is noted in its consideration for both vagueness and intensity of multidimensional deprivation and produces a complete set of synthetic indicators for the description of deprivation on statistical populations. In addition, it also allows for the introduction of information on attribute importance and it facilitates the handling of missing data.
  • In Fattore[129] the procedure “fuzzy-first order dominance”(F-FOD), employs concepts and tools from partially ordered set theory and from fuzzy relational calculus with the goal of overcoming the main limitations of previously developed algorithms for FOD analysis. F-FOD is useful in that it produces full pairwise comparison matrices which allows for partial orderings and rankings of the statistical units to be derived from the input data.
  • Bruggemann et al.[51] also proposes the average rank method which is based on the structure of the Hasse duagram associated to the order relations encoded in the data matrix.